### v1

@@ 3 Neural Networks are Virtual Extended Formulations
We prove this theorem in three steps, represented by the next three propositions. The first proposition shows that every maxout network can be written as the difference of two monotone maxout networks each of which has the same size as the original network. This is related to writing functions represented by neural networks as tropical rational functions [44, 6].
Proposition 1.
The second proposition shows that small monotone neural networks imply small extended formulations of the epigraph. For a convex CPWL function, the epigraph is the polyhedron $\operatorname{epi}(f)=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}\mid t\geq f(x)\}$. The proposition is related to earlier works connecting the extension complexity to Boolean and arithmetic circuit complexity, see e.g. [13, 24].
Proposition 2.
If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is represented by a monotone rank-$k$ maxout network of size $s\geq 1$, then $\operatorname{xc}(\operatorname{epi}(f))\leq ks$.
@@
Based on the neural network, we construct an extended formulation for $\operatorname{epi}(f)$ of size $ks$. For each neuron $v$, we introduce a variable $y_{v}$. For ease of notation, we identify $y_{v}$ for an input neuron $v$ with the corresponding entry of the input vector $x$. Similarly, if $v$ is the output neuron, we identify $y_{v}$ with $t$. Then, the extended formulation for the epigraph is given by the system
@@
Indeed, if $(x,t)\in\operatorname{epi}(f)$, setting $y_{v}\coloneqq z_{v}(x)$ to the value represented by each hidden neuron $v$ yields by construction a feasible solution with respect to all inequalities. Conversely, given a feasible solution $(x,t,y)$, one can inductively propagate the validity of the inequality $y_{v}\geq z_{v}(x)$ along a topological order of the neurons. In particular, for the output neuron, we obtain $t\geq f(x)$ and thus $(x,t)\in\operatorname{epi}(f)$. ∎
Remark 3.
@@
The final ingredient for proving Theorem 3.1 is that the extension complexity of a polytope equals the extension complexity of the epigraph of its support function. This is related to other previous results about the extension complexity being (almost) preserved under various notions of duality, see, e.g., [28, 18, 43].
Proposition 3.
@@
The statement simply follows from combining Propositions 1, 2 and 3. The factor $4$ arises from one factor $2$ through Proposition 1 and another factor $2$ from Proposition 2 as $\operatorname{nnc}(P)$ is defined with $k=2$. ∎
4 Implications for Monotone Neural Networks
The ultimate goal of connecting neural networks and extended formulations in this paper is to use existing powerful results for lower bounds on the extension complexity to obtain lower bounds on neural networks. To pursue this for general neural networks though, we would need lower bounds on the (potentially stronger) notion of virtual extension complexity, which are currently unavailable. However, if we restrict our attention to monotone neural networks, lower bounds on the extension complexity immediately translate to lower bounds on this subclass of neural networks, as we demonstrate.
@@ 4 Implications for Monotone Neural Networks
Combine Propositions 2 and 3. The factor $2$ arises from Proposition 2 as $\operatorname{mnnc}(P)$ is defined with $k=2$. ∎
In particular, we obtain the following bounds on monotone neural networks based on known bounds on the extension complexity. For a complete graph $(V,E)$ with $n$ vertices and a weight vector $c=(c_{e})_{e\in E}\in\mathbb{R}^{E}$, let $f_{\mathrm{MAT}}(c)$ be the value of a maximum weight matching and $f_{\mathrm{TSP}}(c)$ the value of a longest [^1] This is to make the function convex; this is equivalent to finding the shortest tour with respect to the negated weights. traveling salesperson (TSP) tour with respect to $c$. We denote the matching polytope by $P_{\mathrm{MAT}}$ and the TSP polytope by $P_{\mathrm{TSP}}$.
@@
If a monotone maxout network represents $f_{\mathrm{MAT}}$ or $f_{\mathrm{TSP}}$, then it must have size at least $2^{\Omega(n)}$.
Proof.
The functions $f_{\mathrm{MAT}}$ and $f_{\mathrm{TSP}}$ are the support functions of $P_{\mathrm{MAT}}$ and $P_{\mathrm{TSP}}$, respectively. Both have extension complexity $2^{\Omega(n)}$ [37]. Now, the result follows from Theorem 4.1. ∎
In the same way, one can prove lower bounds for neural networks computing the support function of any polytope with high extension complexity, e.g., for neural networks solving the MAX-CUT problem or the stable set problem. The corresponding lower bounds on the extension complexity were first proven by Fiorini, Massar, Pokutta, Tiwary, and de Wolf [14]. Furthermore, they also derived that the extension complexity of the TSP polytope is at least $2^{\Omega(\sqrt{n})}$, before Rothvoß [37] improved it to $2^{\Omega(n)}$.
Theorem 4.3 ([36]).
In fact, the proof of this theorem in [36] shows that almost all matroid base polytopes must have exponential extension complexity. With Theorem 4.1, we get the following implication. For a family of matroids $M_{n}$ on $n$ elements and a weight vector $c\in\mathbb{R}^{n}$, let $f_{n}(c)$ be the value of the maximum weight basis in $M_{n}$.
@@
There exists a family of matroids $M_{n}$ on $n$ elements such that every monotone maxout neural network representing $f_{n}$ must have size $2^{\Omega(n)}$.
For machine learning applications, it is arguably less important to obtain neural networks that exactly represent a given function. Instead, approximating the desired output is often sufficient. We demonstrate that also in the approximate setting, lower bounds on the extension complexity can be transferred to neural networks. We would like to emphasize that there exist many different ways to define what it means to approximate a given function or problem “sufficiently well”, both in optimization and machine learning. The “correct” notion always depends on the context. See also [30] for a discussion of how approximations can be translated between polytopes and neural networks.
@@
Suppose a monotone maxout neural network represents a function $f$ such that $f_{\mathrm{MAT}}(c)\leq f(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. Then the neural network must have size at least $2^{\Omega(\min\{n,1/\epsilon\})}$.
In particular, monotone neural networks cannot serve as fully-polynomial approximation schemes for the matching problem, in the sense that they cannot have polynomial size in both $n$ and $1/\epsilon$ to approximate the matching problem to $\epsilon$-precision. We understand Theorem 4.5 as a prototype for the fact that, in principle, inapproximability results from extended formulations can be transferred to monotone neural networks. See, e.g., a framework for such lower bounds in [7] and inapproximability up to a factor $2$ for vertex cover in [2].
5 Properties of Virtual Extension Complexity
In this section, we give some insights towards understanding the concept of virtual extension complexity. First, we argue that, even though virtual extended formulations are more general than the ordinary extended formulations, they can still be used to optimize efficiently via linear programming. In the second part of the section, we give an example demonstrating that Minkowski sum can indeed drastically reduce the extension complexity. More precisely, there exist polytopes $P+Q=R$ with $\operatorname{xc}(R)$ being way smaller than $\operatorname{xc}(P)$. This means that for obtaining useful lower bounds on $\operatorname{vxc}(P)$, we indeed need to look at both, $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$, and cannot just focus on $\operatorname{xc}(R)$.
5.1 Optimizing over Virtual Extended Formulations
@@
The proof is based on the following fact about faces of Minkowski sums, which can be found, e.g., in [19].
Lemma 1.
@@
Proposition 4 implies that, for generic objective directions $c$, we can efficiently find the optimal solution within $P$ by linear programming using small extended formulations for $Q$ and $R$. However, the situation becomes more tricky if multiple optimal solutions exist within $R$, and then potentially also within $Q$. Indeed, in this case we would not know which of the optimal solutions we should subtract in order to obtain a solution that is feasible for $P$. Therefore, we need to be more careful to handle non-generic objective functions $c$. One solution would of course be to randomly perturb the objective function $c$ by a tiny amount such that the optimal solution in $R$ becomes unique. In practice, however, it is not clear how to ensure that the perturbation is actually small enough and generic enough at the same time.
@@
5.2 Extension Complexity and Minkowski Sums
@@
$\Pi_{n}=\operatorname{conv}\{(1,2,\ldots,n)^{\sigma}\mid\sigma\in S_{n}\}$
arising as the convex hull of all permutations of the vector $(1,2,\ldots,n)$. The regular permutahedron has the property that all edge directions are of the form $e_{i}-e_{j}$ for canonical unit vectors $e_{i},e_{j}$. More generally, polytopes fulfilling this form the class of generalized permutahedra. A different characterization is in terms of Minkowski sums: a polytope $P$ is a generalized permutahedron if and only if there is a polytope $Q$ with $P+Q=\lambda\cdot\Pi_{n}$ for some $\lambda\geq 0$, see [34, Proposition 3.2]. It turns out that every matroid base polytope is such a generalized permutahedron, and we can even choose $\lambda=1$ in this case [4, 15].
@@
For $n\in\mathbb{N}$, there is a polytope $\Pi_{n}$ with extension complexity $\mathcal{O}(\text{poly}(n))$, such that a Minkowski summand has extension complexity $2^{\Omega(n)}$.
@@
By [16], the extension complexity of $\Pi_{n}$ is $\Theta(n\log(n))$. As discussed, each matroid base polytope $P$ is a Minkowski summand of $\Pi_{n}$. Now, Theorem 4.3 guarantees that there exist such polytopes with extension complexity $2^{\Omega(n)}$. ∎
@@
6 Open Questions
Clearly, the most intriguing open question is: Can $\operatorname{vxc}(P)$ be (much) smaller than $\operatorname{xc}(P)$? If the answer is no, then this would imply strong lower bounds on neural networks. On the other hand, if the answer is yes, it implies that solving two linear programs and taking the difference is (much) more powerful than just solving one linear program. In fact, we are not even aware of an example for which we can prove that $\operatorname{vxc}(P)<\operatorname{xc}(P)$. On the other hand, as for the usual extension complexity, we do not expect $\operatorname{vxc}(P)$ to be polynomial for any polytope $P$ associated with an NP-hard optimization problem. An obvious, but seemingly challenging approach to lower-bound virtual extension complexity is to take any previously successful method for lower-bounding extension complexity and try to adapt it such that it also works in the virtual case.
We hope that this paper inspires researchers to answer this question: Are two LPs better than one?
@@ Acknowledgments
Christoph Hertrich is supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No 101153187—NeurExCo.
Appendix 0.A Omitted proofs
Proof of Proposition 1.
$z_{v}$ $=\max\Big{\{}\sum_{u\in\delta^{\mathrm{in}}_{v}}\big{(}a_{uv}^{i}+\sum_{j\neq i}b_{uv}^{j}\big{)}z_{u}\Bigm{|}i=1,\dots,k\Big{\}}-\sum_{u\in\delta^{\mathrm{in}}_{v}}\sum_{j=1}^{k}b_{uv}^{j}z_{u}\enspace.$ (1)
Proof of Proposition 3.
The key idea behind this statement is the well-known fact that the dual of the epigraph of $f_{P}$ as a cone in $\mathbb{R}^{d+1}$ equals the homogenization of $P$, see, e.g., [3, 35]. We give a short geometric proof of this fact to show the ingredients of the construction. A point $(c,t)$ in the epigraph $\operatorname{epi}(f_{P})$ fulfills the equivalent conditions

### v2

@@ 3. Lower Bounds for Monotone and Input-Convex Neural Networks
Before we dive into proving Theorem 3.1, however, we argue that for computing a monotone function, monotone and input-convex maxout networks are equally efficient.
Proposition 3.2.
If a maxout ICNN represents a monotone function $f$, then it can be converted into a monotone neural network of the same size that represents the same function.
Observe that the support function $f_{P}$ of a polytope $P$ is monotone if and only if $P$ is contained in the nonnegative orthant $\mathbb{R}_{\geq 0}^{d}$. Proposition 3.2 implies that for the definition of $\operatorname{mnnc}(P)$ of a polytope $P\in\mathbb{R}_{\geq 0}^{d}$, it does not matter whether we use monotone or input-convex neural networks.
We now prove Theorem 3.1 in two steps, represented by the next two propositions. The first proposition shows that small ICNNs imply small extended formulations of the epigraph. For a convex CPWL function, the epigraph is the polyhedron $\operatorname{epi}(f)=\{(x,t)\in\mathbb{R}^{d}\times\mathbb{R}\mid t\geq f(x)\}$. The proposition is related to earlier works connecting the extension complexity to Boolean and arithmetic circuit complexity, see e.g., Fiorini et al. (2021); Hrubeš and Yehudayoff (2023). A similar statement also appears in the paper introducing ICNNs, where Amos et al. (2017) prove that ICNN inference can be written as a linear program.
Proposition 3.3.
If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is represented by a rank-$k$ maxout ICNN of size $s\geq 1$, then $\operatorname{xc}(\operatorname{epi}(f))\leq ks$.
@@
Based on the ICNN, we construct an extended formulation for $\operatorname{epi}(f)$ of size $ks$. For each neuron $v$, we introduce a variable $y_{v}$. For ease of notation, we identify $y_{v}$ for an input neuron $v$ with the corresponding entry of the input vector $x$. Similarly, if $v$ is the output neuron, we identify $y_{v}$ with $t$. Then, the extended formulation for the epigraph is given by the system
@@
Indeed, if $(x,t)\in\operatorname{epi}(f)$, setting $y_{v}\coloneqq z_{v}(x)$ to the value represented by each hidden neuron $v$ yields by construction a feasible solution with respect to all inequalities. Conversely, given a feasible solution $(x,t,y)$, one can inductively propagate the validity of the inequality $y_{v}\geq z_{v}(x)$ along a topological order of the neurons. Note that, here in the induction step, we need to use that weights between hidden neurons are nonnegative, as the network is input-convex. The potentially negative weights on outgoing edges of the input neurons are no problem because no other variable is multiplied by them in the propagation. Then, for the output neuron, we obtain $t\geq f(x)$ and thus $(x,t)\in\operatorname{epi}(f)$. ∎
Remark 3.4.
@@
The second ingredient for proving Theorem 3.1 is that the extension complexity of a polytope equals the extension complexity of the epigraph of its support function. This is related to other previous results about the extension complexity being (almost) preserved under various notions of duality, see, e.g., Martin (1991); Gouveia et al. (2013); Weltge (2015).
Proposition 3.5.
@@
Combine Propositions 3.3 and 3.5. The factor $2$ arises from Proposition 3.3 as $\operatorname{mnnc}(P)$ is defined with $k=2$. ∎
Now we apply Theorem 3.1 to obtain strong lower bounds on monotone and input-convex neural networks based on known bounds on the extension complexity. For a complete graph $(V,E)$ with $n$ vertices and a weight vector $c=(c_{e})_{e\in E}\in\mathbb{R}^{E}$, let $f_{\mathrm{MAT}}(c)$ be the value of a maximum weight matching and $f_{\mathrm{TSP}}(c)$ the value of a longest [^3] This is to make the function convex; this is equivalent to finding the shortest tour with respect to the negated weights. traveling salesperson (TSP) tour with respect to $c$. We denote the matching polytope by $P_{\mathrm{MAT}}$ and the TSP polytope by $P_{\mathrm{TSP}}$.
@@
If a monotone or input-convex maxout network represents $f_{\mathrm{MAT}}$ or $f_{\mathrm{TSP}}$, then it must have size at least $2^{\Omega(n)}$.
@@
The functions $f_{\mathrm{MAT}}$ and $f_{\mathrm{TSP}}$ are the support functions of $P_{\mathrm{MAT}}$ and $P_{\mathrm{TSP}}$, respectively. Both have extension complexity $2^{\Omega(n)}$ (Rothvoß, 2017). Now, the result follows by Theorem 3.1. ∎
In the same way, one can prove lower bounds for neural networks computing the support function of any polytope with high extension complexity, e.g., for neural networks solving the MAX-CUT problem or the stable set problem. The corresponding lower bounds on the extension complexity were first proven by Fiorini et al. (2015), who also derived that the extension complexity of the TSP polytope is at least $2^{\Omega(\sqrt{n})}$, before Rothvoß (2017) improved it to $2^{\Omega(n)}$.
Theorem 3.7 (Rothvoß (2013)).
In fact, the proof of this theorem in Rothvoß (2013) shows that almost all matroid base polytopes must have exponential extension complexity. With Theorem 3.1 and Proposition 3.2, we get the following implication. For a family of matroids $M_{n}$ on $n$ elements and a weight vector $c\in\mathbb{R}^{n}$, let $f_{n}(c)$ be the value of the maximum weight basis in $M_{n}$.
@@
There exists a family of matroids $M_{n}$ on $n$ elements such that every monotone or input-convex maxout neural network representing $f_{n}$ must have size $2^{\Omega(n)}$.
For machine learning applications, it is arguably less important to obtain neural networks that exactly represent a given function. Instead, approximating the desired output is often sufficient. We demonstrate that also in the approximate setting, lower bounds on the extension complexity can be transferred to neural networks. We would like to emphasize that there exist many different ways to define what it means to approximate a given function or problem “sufficiently well”, both in optimization and machine learning. The “correct” notion always depends on the context. See also Misiakos et al. (2022) for a discussion of how approximations can be translated between polytopes and neural networks.
@@
Suppose a monotone or input-convex maxout neural network represents a function $f$ such that $f_{\mathrm{MAT}}(c)\leq f(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. Then the neural network must have size at least $2^{\Omega(\min\{n,1/\epsilon\})}$.
In particular, monotone and input-convex neural networks cannot serve as fully-polynomial approximation schemes for the matching problem, in the sense that they cannot have polynomial size in both $n$ and $1/\epsilon$ to approximate the matching problem to $\epsilon$-precision. We understand Theorem 3.9 as a prototype for the fact that, in principle, inapproximability results from extended formulations can be transferred to neural networks. See, e.g., a framework for such lower bounds in Braun et al. (2015) and inapproximability up to a factor $2$ for vertex cover in Bazzi et al. (2019).
4. Virtual Extension Complexity
In this section, we study the novel concept of virtual extension complexity $\operatorname{vxc}(P)$. We first prove that $\operatorname{vxc}(P)$ lower-bounds the size of general neural networks, motivating the study of $\operatorname{vxc}(P)$ from a machine learning perspective. We then argue that, even though virtual extended formulations are more general than the ordinary extended formulations, they can still be used to optimize efficiently via linear programming, motivating the study of $\operatorname{vxc}(P)$ from a combinatorial optimization perspective. Finally, we give an example demonstrating that Minkowski sum can indeed drastically reduce the extension complexity. More precisely, there exist polytopes $P+Q=R$ with $\operatorname{xc}(R)$ being way smaller than $\operatorname{xc}(P)$. This means that for obtaining useful lower bounds on $\operatorname{vxc}(P)$, we indeed need to look at both, $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$, and cannot just focus on $\operatorname{xc}(R)$.
4.1. Neural Networks are Virtual Extended Formulations
We prove the following theorem stating that neural network sizes can be lower-bounded through virtual extension complexity.
@@ 4.1. Neural Networks are Virtual Extended Formulations
The proof of this theorem requires three steps, two of which are already available through Propositions 3.3 and 3.5. In addition, we need another proposition showing that every maxout network can be written as the difference of two monotone maxout networks each of which has the same size as the original network. This is related to writing functions represented by neural networks as tropical rational functions (Zhang et al., 2018; Brandenburg et al., 2024b).
Proposition 4.2.
Proof of Theorem 4.1..
The statement simply follows from combining Propositions 4.2, 3.3 and 3.5. The factor $4$ arises from one factor $2$ through Proposition 4.2 and another factor $2$ from Proposition 3.3 as $\operatorname{nnc}(P)$ is defined with $k=2$. ∎
4.2. Optimizing over Virtual Extended Formulations
@@
The proof is based on the following fact about faces of Minkowski sums, which can be found, e.g., in Grünbaum (2003).
Lemma 4.4.
@@
Proof of Proposition 4.3..
As for $R$ the optimal face in $c$-direction is the singleton $\{x^{R}\}$, by Lemma 4.4 the corresponding faces of $P$ and $Q$ must be singletons, too, namely $\{x^{P}\}$ and $\{x^{Q}\}$ with $x^{P}+x^{Q}=x^{R}$. ∎
Proposition 4.3 implies that, for generic objective directions $c$, we can efficiently find the optimal solution within $P$ by linear programming using small extended formulations for $Q$ and $R$. However, the situation becomes more tricky if multiple optimal solutions exist within $R$, and then potentially also within $Q$. Indeed, in this case we would not know which of the optimal solutions we should subtract in order to obtain a solution that is feasible for $P$. Therefore, we need to be more careful to handle non-generic objective functions $c$. One solution would of course be to randomly perturb the objective function $c$ by a tiny amount such that the optimal solution in $R$ becomes unique. In practice, however, it is not clear how to ensure that the perturbation is actually small enough and generic enough at the same time.
@@
4.3. Extension Complexity and Minkowski Sums
@@
$\Pi_{n}=\operatorname{conv}\{(\sigma(1),\sigma(2),\ldots,\sigma(n))\mid\sigma\text{ is a permutation of }\{1,\dots,n\}\}$
arising as the convex hull of all permutations of the vector $(1,2,\ldots,n)$. The regular permutahedron has the property that all edge directions are of the form $e_{i}-e_{j}$ for two unit vectors $e_{i},e_{j}$. More generally, polytopes fulfilling this form the class of generalized permutahedra. A different characterization is in terms of Minkowski sums: a polytope $P$ is a generalized permutahedron if and only if there is a polytope $Q$ with $P+Q=\lambda\cdot\Pi_{n}$ for some $\lambda\geq 0$, see Postnikov et al. (2008, Proposition 3.2). It turns out that every matroid base polytope is such a generalized permutahedron, and we can even choose $\lambda=1$ in this case (Borovik et al., 2003; Fujishige, 2005).
@@
For $n\in\mathbb{N}$, there is a polytope $\Pi_{n}$ with extension complexity $\mathcal{O}(\text{poly}(n))$, such that a Minkowski summand of $\Pi_{n}$ has extension complexity $2^{\Omega(n)}$.
@@
By Goemans (2015), the extension complexity of the regular permutahedron $\Pi_{n}$ is of the order $\Theta(n\log(n))$. As discussed, each matroid base polytope $P$ is a Minkowski summand of $\Pi_{n}$. Now, Theorem 3.7 guarantees that there exist such polytopes with extension complexity $2^{\Omega(n)}$. ∎
@@
5. Conclusions
In this paper we focused on proving strong lower bounds for monotone and input-convex neural networks based on existing breakthrough results about extension complexity. Furthermore, we paved the way towards transferring these bounds to general neural networks by introducing the notion of virtual extension complexity and relating it to neural networks.
An obvious question for future research is a thorough analysis of how much more powerful general neural networks are compared to monotone or input-convex neural networks. Can we find a function for which there is an exponential gap in the required size?
Closely related, and from the perspective of polyhedral theory and combinatorial optimization, the most intriguing open question is: Can $\operatorname{vxc}(P)$ be (much) smaller than $\operatorname{xc}(P)$? If the answer is no, then this would imply strong lower bounds on neural networks. On the other hand, if the answer is yes, it implies that solving two linear programs and taking the difference is (much) more powerful than just solving one linear program. In fact, we are not even aware of an example for which we can prove that $\operatorname{vxc}(P)<\operatorname{xc}(P)$. On the other hand, as for the usual extension complexity, we do not expect $\operatorname{vxc}(P)$ to be polynomial for any polytope $P$ associated with an NP-hard optimization problem. An obvious, but seemingly challenging approach to lower-bound virtual extension complexity is to take any previously successful method for lower-bounding extension complexity and try to adapt it such that it also works in the virtual case.
To summarize, we hope that this paper inspires researchers to find out: Are two LPs better than one?
@@ Acknowledgments
Part of this work was completed while Christoph Hertrich was affiliated with Université Libre de Bruxelles, Belgium, and received support by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No 101153187—NeurExCo.
Appendix A Omitted Proofs
Proof of Proposition 3.2.
Let $\mathcal{N}_{0}$ be the maxout ICNN of size $s$ representing $f$. Without loss of generality we assume that every input neuron has a connection to every maxout unit. We can always achieve this by introducing arcs with weights equal to zero. This does not increase the number of neurons and therefore not the size of the neural network. Let $v_{1},v_{2},\dots,v_{s}$ be a topological order of the maxout units of $\mathcal{N}_{0}$.
We construct a sequence of equivalent neural networks $\mathcal{N}_{1},\dots,\mathcal{N}_{s-1}$, removing incoming negative weights neuron by neuron. We only change weights on arcs leaving input neurons. All other weights remain unchanged in this process, as they are already nonnegative, because $\mathcal{N}_{0}$ is input-convex. Our construction ensures that in $\mathcal{N}_{p}$ none of the neurons $v_{1}$ to $v_{p}$ will have negative weights on their incoming arcs.
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first $p-1$ maxout neurons unchanged. Let $I\subseteq V$ be the set of input neurons. Consider the weights $w_{uv}^{i}$ on arcs entering $v\coloneqq v_{p}$ from an input neuron $u\in I$. Let $\gamma_{u}\coloneqq\min_{i=1,\dots,k}w_{uv}^{i}$ be the smallest such weight for each $u\in I$, which might be negative or positive. Observe that
(1) $z_{v}$ $=\max\Big{\{}\sum_{u\in I}(w_{uv}^{i}-\gamma_{u})z_{u}+\sum_{u\in\delta^{\mathrm{in}}_{v}\setminus I}w_{uv}^{i}z_{u}\Bigm{|}i=1,\dots,k\Big{\}}+\sum_{u\in I}\gamma_{u}z_{u}\enspace.$
This allows to construct $\mathcal{N}_{p}$ from $\mathcal{N}_{p-1}$ by the following operation. Each weight $w_{uv}^{i}$ of $v=v_{p}$ coming from an input neuron $u\in I$ is set to $w_{uv}^{i}-\gamma_{u}\geq 0$. To make up for this change, one has to correct the weights from input neurons to neurons $v_{q}$ with $q>p$ to incorporate the term outside the maximum in (1). More precisely, if $\widetilde{v}\coloneqq v_{q}$ is an out-neighbor of $v=v_{p}$ and $u\in I$ is an input neuron, then we need to update the weight $w_{u\widetilde{v}}^{i}$ to $w_{u\widetilde{v}}^{i}+w_{v\widetilde{v}}^{i}\gamma_{u}$. It is easy to verify that this weight update exactly recovers the missing term from (1) when computing $z_{\widetilde{v}}$. This might introduce new negative weights, but all of them are associated with arcs from input neurons to neurons in the topological order after $p$. So, after doing this operation, we obtain an equivalent neural network $\mathcal{N}_{p}$ with only nonnegative weights associated with the neurons $v_{1}$ to $v_{p}$.
We continue that way until we obtain $\mathcal{N}_{s-1}$, where the only weights that could potentially be negative are those between an input neuron and the output neuron $v_{s}$. We show that, in fact, these weights are nonnegative, too, and $\mathcal{N}_{s-1}$ is our desired monotone neural network computing $f$.
To this end, we first show by induction on $p=1,\dots,s-1$ that for a negative unit vector $x=-e_{j}$ as input, each neuron $v=v_{p}$ outputs $z_{v}(-e_{j})=0$. If $p=1$, then the only incoming edges of $v=v_{1}$ are coming from input neurons. Moreover, since we are inputting $-e_{j}$, only the $j$-th input neuron, call it $u$, propagates a nonzero value. By the choice of $\gamma_{u}$, there must be an index $i$ such that our updated weight $w_{uv}^{i}-\gamma_{u}$ is exactly zero, while it is nonnegative for all other indices. Therefore, upon input $-e_{j}$, the maximum expression evaluates to 00, settling the induction start. For the induction step, observe that the very same argument applies, since by induction all previous neurons output zero upon input $-e_{j}$.
As a consequence of the latter claim, we obtain that the output $z_{v_{s}}(-e_{j})$ of $\mathcal{N}_{s-1}$ is precisely the negated weight from $u$ to $v_{s}$. Since $z_{v_{s}}(0)=0$, monotonicity of the final function implies that the output on $-e_{j}$ must be nonpositive, implying that the weight must be nonnegative, as claimed. Thus, we just showed that also the incoming weights of $v_{s}$ must be nonnegative in $\mathcal{N}_{s-1}$, implying that $\mathcal{N}_{s-1}$ is monotone. ∎
Proof of Proposition 3.5..
The key idea behind this statement is the well-known fact that the dual of the epigraph of $f_{P}$ as a cone in $\mathbb{R}^{d+1}$ equals the homogenization of $P$, see, e.g., Bertsekas [2009], Rockafellar [1970]. We give a short geometric proof of this fact to show the ingredients of the construction. A point $(c,t)$ in the epigraph $\operatorname{epi}(f_{P})$ fulfills the equivalent conditions